Thursday night, Jan. 6, 2011

Observers: Jill Bechtold, Dave Thompson

Afternoon: Completely cloudy.

Sunset: Fairly cloudy but will try Lucifer imaging.

5pm. Switching from LBC to Lucifer

6pm. Can see the moon, but very cloudy.

6:30pm. Problems with writing to the archive, Dave is on the phone with Steven to fix.

6:30pm. Very cloudy, but opening. Operator is going to pointing star.

6:45pm Telescope won't go to a commanded ra,dec and track. John Little here to troubleshoot. Encoder problem?

6:53pm. Archive file writing problem fixed!

6:58pm. On phone to servo engineer. Aximuth encoder problem. "Friction too high." Overshoots commanded position by degrees and oscillates. "Rails"

7:00 pm increased oil pressure on az axis pads. looking at telemetry. Calling more people.

Dave says to type rm clouds

7:20pm Operator off to do major restart of computer "upstairs".

8:09pm. Operator returns, finally. There's a major software problem with "JET" the computer which controls the telescope mount servos. There was a power outage this morning, which may have something to do with it. Dave Ashby and Dan Cox are working on it.

Clouds are thinning, but it's still cloudy.

8:30pm John Little is going "upstairs" to try to bypass broken parts with a usb cable, or swap boards, with Dan Cox and Dave Ashby on the phone.

9pm. Computer hardware swap was successful. Kernel is being reloaded.

9:20pm. John Little says "problem solved". Turns out none of this JET stuff was the problem. Someone had not removed CHOCKS that were impeding the telescope progress. These are supposed to removed during the "Walk Through" before the beginning of the night. It's on the checklist, but was not checked off by the person filling the dewars, etc. Sigh...

9:30pm. Going to pointing star to focus and collimate. Seeing 1.2-1.5 arcsec.

9:40pm Observing MACS 0744 Ks backup Herschel Cluster. Changed the dither strategy, one hour exposure.

DIT=10, NDIT=6, jitter 60 30

also changed SAVE_MODE = INTEGRATED instead of normal -- we don't want 6 files at every dither position.

Didn't want to spend 2 min at each dither point in this weather

Looking good! Except it's still cloudy and seeing is 1.2-1.3 arcsec

10:09pm after about 15 min of exposure, we've lost star in shack-hartmann, still guiding, sort of. Lots of clouds.

10:45pm Things are a lot better. seeing .7". still cloudy but some of the image differences are really good

11:29pm Started another 20 dithers on MACS 0744. had to update pointing because we couldn't just start where we ended. so next time will comment out Telescope_Setup commands in script if just continuing on object. This will be fixed when they fix the pointing models in Lucifer.

So it took about 2 hours of real time to get 60 minutes of open shutter exposure.

From the way the sky looks, it's a miracle we're getting anything at all, but it's not half bad.

11:53pm It's getting cloudy. Last images of this dither sequence are not very good.

12:02am Moving to Ian's High z Lens, J0946+1835 in z

did J0946+1825_zacq.txt, then J0946+1825_z.txt

12:11am Now doing J. did _Jacq.txt, then _J.txt script twice. Clouds, but seeing ok.

12:20am Doing H. and K. Very good

12:38am Going to McGreer's radio loud quasar 1054+4541

had to change dither offset type to absolute doing H first CLOUDS

Guide star too dim, failed.

looked up guide stars, this one has R=16, it's too faint.

punted, but maybe the problem was with the collimation, see below

12:52am Off to Abell 0611 for Marie. Edited parameters again. Will do 30 minutes here

CLOUDS. script stopped, can't see guide star

A bit of futzing around -- collimation at limits, needed to reset and repoint

started exposure at 1:19am Seeing .4" Very cool!

2:08am Going to MACS1149. First operator will do pointing and collimation.

2:14 started exposing -- macs1149.K.txt 30 minute exposure

Guiding etc but it's very cloudy.

Lost guide star, script stopped. no good data obtained so far.

2:32am Closed. Too cloudy.

Opaque at MMT also.
Topic revision: r5 - 12 Mar 2011, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback